OpenStudy (anonymous):

WIN MEDAL! Which ordered pairs are solutions to the equation? y = x – 4 A. (10, 6) and (5, 1) B. (10, 5) and (5, 1) C. (10, 4) and (5, 1) D. (10, 6) and (1, 5)

OpenStudy (anonymous):

WIN MEDAL FOR BEST ANSWER!!

OpenStudy (anonymous):

I do not know how to do it... i need an answer.

OpenStudy (anonymous):

lol just put the values its A

OpenStudy (anonymous):

Thank you! @theblademaster

OpenStudy (anonymous):

y= 10-4 = 6 y= 5-1 =1 so 10,6 and 5,1 is justified

OpenStudy (neer2890):

if you got any point (x1,y1) for any line then it is obvious that it satisfies that line. put x1 inplace of x and y1 in place of y. if it satisfies the equation then it's correct.

OpenStudy (anonymous):

@aajugdar Win Medal

OpenStudy (anonymous):

5-4*

OpenStudy (uri):

You're not getting an answer. WE can only guide you to the answer.Giving direct answer is restricted. Please read the Coc : http://openstudy.com/terms-and-conditions

OpenStudy (anonymous):

lol this is a basic question you dont need to go throu procedure for it? and lol i can tell who wants the explanation and who doesnt half of them come to cheat here in online exam lets not make this a thread ^^ ta-tah